OpenStudy (anonymous):

Hockey teams receive 2 points for a win and 1 point for a tie. The Wildcats once won a championship with 60 points. They won 9 more games than they tied. How many wins and how many ties did the Wildcats have?

OpenStudy (jlastino):

x= games tied x+9= games won x+ 2(x+9)= 60 x+2x+18=60 3x=60-18 3x=42 x=14 - ties x+9=23 - games won

OpenStudy (anonymous):

23 wins and 14 ties

OpenStudy (jlastino):

Yeah

OpenStudy (anonymous):

thanks